If you earn 6% per year on your bank account, how long will it take an account with $110 to double to $220? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

FV = PV × (1 + r)t

$220 = $110 × 1.06t

1.06t = 2

t × ln1.06 = ln2

t = ln2 / ln1.06

t = 11.90 years